Healty people have body temperatures that are normally distributed with a mean of 98.20∘F and a standard deviation of 0.62∘F .
(a)  If a healthy person is randomly selected, what is the probability that he or she has a temperature above 100∘F?
(b)  A hospital wants to select a minimum temperature for requiring further medical tests. What should that temperature be, if we want only 0.5 % of healty people to exceed it?
Answer #1

solution Given that el = 98.20 6 = 0.62 o) P(x>100) = 1-pc x<100) = IP/x-W 100-98.20 ( 6 0.62 = 1- p c 1.8 ze 0,62 = 1- p (2xb) using standard normal table p (272) : 0.57. I-p(222) = 0,005 p(222) = 1-0.005 - 0.995 p(Z< 2.58 ) = 0,995 z = 2.58 using a